# SVLA: A Unified Speech-Vision-Language Assistant with Multimodal Reasoning and Speech Generation

### v1

@@ Abstract
Large Vision-Language Models have shown impressive capabilities in tasks such as image captioning, visual question answering, and cross-modal retrieval. However, there are still significant challenges that need to be addressed in order to fully unlock the potential of these models. First, integrating speech, text, and vision into a unified model is particularly difficult for tasks like Spoken Image Captioning and Spoken Visual Question Answering, where the interaction between these modalities introduces additional complexity. Second, existing speech generation approaches differ—some generate speech directly, while others use an intermediate text step—but their impact on fluency, coherence, and accuracy remains unexplored. Third, current speech evaluation relies heavily on Automatic Speech Recognition, which introduces biases and fails to directly assess speech quality. To address these challenges, we propose SVLA, a unified Speech-Vision-Language Assistant based on a decoder-only transformer architecture that seamlessly integrates multimodal inputs and outputs. We enhance model performance with a large-scale speech-text-image dataset containing 38.2 million examples and 64.1 hours of TTS-generated speech. Additionally, we introduce Speech-VQA Accuracy, a novel acoustic evaluation metric for assessing speech-based VQA tasks. Our approach advances multimodal understanding and generation, facilitating more effective integration of speech, text, and vision (http://github.com/vlm-svla/svla).
@@ 1 Introduction
Large Language Models (LLMs) like LLaMA [1], GPT-3 [2], InternLM-2 [3], and Qwen [4] excel in textual processing but are limited by their unimodal design. To address real-world needs, Large Vision-Language Models (LVLMs) such as GPT-4V [5], Gemini [6], BLIP-2 [7], Llava [8], and Flamingo [9] integrate visual data, while Unified Speech-Text-Language Models further incorporate speech to enhance accessibility and natural interactions. Recent efforts—exemplified by closed-source ChatGPT-4o [10] and open-source models like LLaMA 3.1 [11], AnyGPT [12], and NextGPT [13]—have expanded multimodal capabilities to speech-vision tasks, such as generating images from speech or creating music from visual inputs.
However, despite the promising advancements made so far, significant challenges remain that must be addressed in order to fully unlock the potential of these models. First, these models lack robust reasoning abilities for tasks like image captioning (IC) and visual question answering (VQA), where inputs may be text or speech and outputs may be speech or text, limiting their applicability to complex multimodal interactions. Second, existing speech generation approaches vary, with some models generating speech directly, while others use text instructions as an intermediary step. However, no prior research has systematically compared these two techniques. This distinction is crucial, as the choice between these paradigms significantly impacts fluency, coherence, and response accuracy in multimodal AI. For example, when asked “What are the people doing?”, a text-instruction model first generates a structured textual response before generating speech tokens (e.g., “The textual answer is ‘running.’ Therefore, the speech answer is ‘running.’(speaking)”). While this ensures consistency between text and speech outputs, it introduces redundancy and unnatural phrasing, which may not align with natural spoken communication. In contrast, direct speech generation bypasses the textual constraint, producing more fluid spoken responses but at the risk of greater variability, pronunciation inconsistencies, or a loss of linguistic structure. Understanding these trade-offs is essential for advancing multimodal reasoning, spoken language generation, and user interaction quality in AI systems. Finally, existing evaluation metrics for multimodal learning focus mainly on text and vision, offering limited insight into speech-based tasks.
To address the challenges mentioned above, we introduce SVLA, a Speech-Vision-Language Assistant—a unified multimodal system designed for end-to-end reasoning and generation. A speech discretization tokenizer converts raw speech into discrete semantic tokens, enabling the model to process speech similarly to text. This design allows a traditional generative LLM to seamlessly handle tri-modal inputs—speech, text, and images—and generate outputs in either speech or text. SVLA supports a broad range of tasks, including ASR, TTS, Image Captioning (IC), and Visual Question Answering (VQA). Importantly, the system is flexible in handling various modality combinations. For example, it can generate text outputs from speech-based inputs (speech-to-text tasks) or produce speech outputs from text-based inputs (text-to-speech tasks). To evaluate SVLA, we propose Acoustic Speech-VQA Accuracy, a novel metric that directly assesses spoken responses—bypassing ASR conversion—providing a more accurate and unbiased evaluation of the model’s spoken language capabilities. Additionally, SVLA is assessed on a new, large-scale speech-text-vision dataset that empowers models to perform visual tasks using both text and speech. This dataset is built by integrating several widely-used multimodal and speech resources, including VQAv2[14], LAION-600M[15], Visual Genome[16], LibriHeavy[17], LibriSpeech[18], CommonVoice[19], A-OKVQA[20], VizWiz[21], GQA[22], and COCO-Caption[23]. To ensure natural and diverse speech data, we utilize a publicly available TTS system[24] that adjusts accent and speed, enriching linguistic variety and prosody.
@@
  • We introduce SVLA, a unified multimodal assistant designed for end-to-end speech and text reasoning in visual understanding.
  • We present a large-scale speech-text-vision dataset built from widely used multimodal and speech benchmarks, providing aligned text and speech annotations to enhance multimodal learning and facilitate speech-vision-language tasks in both textual and spoken formats.
  • We propose Acoustic Speech-VQA Accuracy, a novel metric for evaluating spoken VQA responses without relying on ASR conversion, ensuring a more direct and unbiased assessment of speech-based reasoning.
  • We establish a comprehensive evaluation framework that systematically benchmarks models across diverse input-output configurations.
@@ 2 Related Works
Below, we explore key advancements in integrating speech into multimodal LLMs, compare direct and instruction-based speech generation approaches, and examine evaluation metrics for speech in multimodal tasks.
Speech Multi-modal Large Language Models: Recent efforts have focused on integrating audio capabilities into LVLMs to enhance multimodal interactions. Early models supported only speech input, limiting their effectiveness in conversational AI [25, 26]. To address this, some approaches incorporated TTS for spoken responses [27, 11], but these pipelines introduce higher computational costs, inference latency, and limited contextual adaptation. OpenAI’s GPT-4o [28] overcomes these limitations by generating speech, enabling real-time, expressive interactions with preserved intonation and emotional cues. However, GPT-4o remains proprietary, prompting open-source efforts to replicate its multimodal capabilities. Models such as Next-GPT [13], Codi-1 [29], Codi-2 [30], UIO [31], and AnyGPT [12] have attempted vision-speech integration using pre-trained encoders. They mostly pay attend on spoken tasks as: Image Generation, Music Generation, Audio Captioning, Video Generation, ASR, and TTS . Yet, most remain limited to IC and VQA within a TTT paradigm, lacking support for crucial speech-vision tasks such as TTS, STT, and STS. However, TMT [32] acttually build a model to speak image captions, but is designed to translate between modalities one at a time, meaning it can handle one input modality at a time and then generate an output in another modality.
Speech-Text-Vision Datasets: While multimodal datasets like VQAv2 [14], LAION-600M [15], and Visual Genome [16] focus on text-vision alignment, they lack natural speech data, limiting their applicability to speech-vision reasoning. Some datasets incorporate audio-visual elements [33, 34, 35, 36], but audio is often restricted to environmental sounds or extracted speech, with questions presented in text. Speech-COCO [37] extends image captioning with speech annotations but remains small in scale and lacks diverse multimodal reasoning tasks. Moreover, existing datasets do not support seamless switching between speech and text, preventing models from dynamically adapting to user intent, such as generating spoken answers or textual captions based on instructions.
Direct vs. Instruction-Based Speech Generation:
Speech generation in multimodal AI follows two main approaches: direct speech generation and instruction-based speech generation. Direct models generate spoken responses end-to-end without intermediate text, while instruction-based models—such as ChatGPT-4o [10], AnyGPT, and Next-GPT [13, 29]—produce both text and speech simultaneously, with text guiding the spoken response. These models output structured text (e.g., ”The answer is ‘running’.”) before generating speech, ensuring consistency but introducing redundancy and unnatural phrasing. While instruction-based generation aligns text and speech outputs, it is less suited for natural conversations. However, a systematic comparison between these approaches remains unexplored, leaving open questions about their effectiveness for context-aware spoken interactions in multimodal AI.
Speech-Based Evaluation Metrics for Multimodal Tasks:
Evaluating text output is straightforward using established metrics like BLEU [38], ROUGE [39], Accuracy, and CIDEr [40]. In contrast, speech evaluation is challenging due to its continuous nature, prosody, and pronunciation variations, which traditional text-based metrics cannot fully capture. A common approach transcribes speech using ASR before applying text metrics [12, 13], but ASR errors can introduce inconsistencies and bias evaluations. Pretrained speech models like WavLM-TDNN [41] have been used to measure similarity in speech output tasks, but they do not account for reasoning quality or semantic correctness. This highlights the need for a dedicated evaluation metric to assessing spoken responses in multimodal reasoning tasks.
@@ 3 Data Generation
A variety of image-text datasets—such as COCO-caption [23], and LAION-COCO [15]—offer abundant and diverse visual and linguistic data. We leverage these resources to build a robust, multi-purpose dataset by pairing each image with its corresponding text and converting that text into speech using the Melon-TTS model [24]. To ensure the generated audio sounds natural, we vary both the speaking speed (ranging from 0.7 to 1.3) and the accents (British, American, Indian, and Australian). Additionally, all audio is standardized at a 16 kHz sampling rate, striking a good balance between quality and size, and maintaining consistent clarity across all recordings.
@@ 3.1 Pre-train dataset
Our pre-training strategy is designed to develop a unified speech-vision-text model capable of handling TTS, ASR, Image Captioning (IC), and VQA. The training data is divided into two key subsets: (1) text-speech data, sourced from LibriHeavy [17], comprising 8M audiobook-transcript pairs to enhance TTS expressiveness and ASR accuracy; and (2) vision-text-speech data, which integrates LAION-COCO [15] (8M image-text pairs) and Visual Genome (VG) [16] to improve the model’s ability to interpret and respond to visual queries. To ensure seamless modality switching, we incorporate instructional prompts such as “Answer this question in speech” or “Write down what you see in the image”. Figure 1 shows the token distribution across three modalities: Image, Text, and Speech [^1] The number of text tokens is computed using Qwen-2.5-1.5B [42], while speech tokens are counted using SpeechTokenizer [43]. Image tokens are estimated at 256 tokens per image.. Additional details on pre-training are provided in Appendix A.
@@
3.2 Visual Instruction dataset
For supervised fine-tuning (SFT), we follow a similar structure but introduce multi-turn conversations to enhance coherence and context retention. The text-speech subset includes LibriSpeech [18] and CommandVoice [19] for diverse linguistic and command-based tasks. The vision-text-speech subset incorporates VQAv2, A-OKVQA [20], GQA [22], VizWiz [21], and COCO-Caption-2014, ensuring broad multimodal coverage. Unlike pre-training, SFT dialogues extend up to 4096 tokens, allowing the model to sustain long-form interactions. Modality-switching prompts guide the model’s transitions between speech, text, and vision tasks, ensuring flexible and adaptive multimodal reasoning. Further details on the SFT dataset can be found in Appendix A.
4 Model Architecture
Our architecture builds on LLaVA’s design [8] with additional components for speech processing. Our model integrates speech and vision inputs into an LLM, denoted as $f_{\theta}(x)$, where $\theta$ represents the model parameters. The overall network architecture is depicted in Figure 2.
@@ 4.3 Multimodal Fusion
Given textual tokens $Z_{t}$ from input text $X_{t}$, discrete speech tokens $Z_{s}$, and visual embeddings $H_{v}$, we construct the multimodal token sequence:
$X=[bos,\,boi,\,img,\,eoi,\,Z_{t},\,boa,\,Z_{s},\,eoa,\,eos],\quad X\in\mathbb{R}^{n}$
where $n$ is the initial token sequence length. Special tokens ($bos,boi,eoi,boa,eoa,eos$) explicitly mark the beginning and end of the sequence, image tokens, and audio tokens.
After embedding, we replace the placeholder embedding of the token $img$ with the visual embeddings $H_{v}$:
$E^{\prime}=[E_{bos},\,E_{boi},\,H_{v},\,E_{eoi},\,E_{Z_{t}},\,E_{boa},\,E_{Z_{s}},\,E_{eoa},\,E_{eos}],$
$\quad E^{\prime}\in\mathbb{R}^{n^{\prime}\times d_{h}}$, where $n^{\prime}=n_{v}+n-1$. Finally, the backbone LLM processes $E^{\prime}$ to achieve multimodal understanding.

### v2

@@ Abstract
Large Vision-Language Models have shown impressive capabilities in tasks such as image captioning, visual question answering, and cross-modal retrieval. However, there are still significant challenges that need to be addressed in order to fully unlock the potential of these models. First, integrating speech, text, and vision into a unified model is particularly difficult for tasks like Spoken Image Captioning and Spoken Visual Question Answering, where the interaction between these modalities introduces additional complexity. Second, existing speech generation approaches differ—some generate speech directly, while others use an intermediate text step—but their impact on fluency, coherence, and accuracy remains unexplored. To address these challenges, we propose SVLA, a unified Speech-Vision-Language Assistant based on a decoder-only transformer architecture that seamlessly integrates multimodal inputs and outputs. We enhance model performance with a large-scale speech-text-image dataset containing 38.2 million examples and 64.1 hours of TTS-generated speech. Our approach advances multimodal understanding and generation, facilitating more effective integration of speech, text, and vision (http://github.com/vlm-svla/svla).
@@ 1 Introduction
Recent advances in Large Vision-Language Models (LVLMs)  [1, 2, 3, 4] mark a significant step toward multimodal AI, enabling models to interpret and reason over visual and textual inputs. However, these models remain constrained by their reliance on text-based instructions and lack native support for speech. Efforts such as LLaMA 3  [5] and Qwen-Audio  [6] have begun incorporating speech through modality-specific encoders, these models remain limited to speech perception and do not support speech generation.
Recent models [7, 8, 9, 10] address this limitation by introducing discrete speech representations, allowing language models to process speech as semantic tokens within a unified speech-text token space. This enables bidirectional modeling and multimodal integration. However, most systems remain focused on shallow tasks—such as text-to-speech [11], image-to-music [12], or basic audio captioning [13]—and struggle with more complex reasoning tasks like image captioning (IC) or visual question answering (VQA) involving spoken inputs or outputs. A key bottleneck is the absence of large-scale, richly aligned datasets spanning speech, text, and vision. Most models are trained on unimodal or bimodal data, limiting their ability to generalize to cognitively demanding, speech-enabled multimodal reasoning.
Furthermore, two prominent paradigms have emerged for enabling speech capabilities within multimodal systems [7, 14]: Cross-modal Instruction and Chain-of-Modality Instruction. Cross-modal Instruction directly maps inputs across modalities, such as converting an image to speech, offering efficiency but frequently compromising semantic coherence. In contrast, Chain-of-Modality Instruction involves first translating the input into textual form, performing reasoning over the generated text, and subsequently producing output in speech or text form. This method leverages the structured reasoning and planning capabilities inherent to large language models but may lead to verbosity. However, a comparative analysis of these paradigms under consistent tri-modal settings within LVLMs has not yet been conducted.
To address the above challenges, we introduce a large-scale synthetic trimodal dataset—combining text, image, and speech—to support instruction-following in both written and spoken formats. This dataset integrates widely-used resources, including VQAv2 [15], LAION-600M [16], Visual Genome [17], LibriHeavy [18], LibriSpeech [19], CommonVoice [20], A-OKVQA [21], VizWiz [22], GQA [23], and COCO-Caption [24]. To ensure natural-sounding and diverse speech data, we use a controllable text-to-speech (TTS) system that varies accent, speaking speed, and prosody. We further include noise and stylistic augmentation from MUSAN [25] to improve robustness. Additionally, we implement a modality-switching instruction tuning strategy, enabling the model to flexibly process and respond in either text or speech.
Building on this foundation, we propose SVLA (Speech-Vision-Language Assistant), a unified, self-supervised multimodal model capable of reasoning over image, text, and speech inputs and generating outputs in either text or speech. SVLA features a hybrid fusion architecture: it applies early fusion between speech and text using discrete semantic units, allowing both to be modeled jointly in a shared language space  [8, 26], and late fusion for visual input by integrating image embeddings from a pretrained encoder  [3]. This design supports complex tasks such as spoken VQA, speech-driven image captioning, and multimodal instruction following.
We also present the systematic evaluation comparing Cross-modal and Chain-of-Modality Instruction within a unified vision-language-speech framework. Our experiments span four controlled configurations—text-to-text, text-to-speech, speech-to-text, and speech-to-speech—using consistent instruction formats and shared example naming across tasks. Benchmarks such as VQAv2 and COCO-Captions are used to assess the trade-offs between fluency, coherence, and semantic quality under each setting.
@@
  • We construct a large-scale, aligned speech-text-vision dataset from diverse benchmarks, supporting instruction-following in both textual and spoken forms.
  • We propose SVLA, a unified tri-modal model that integrates early fusion (text-speech) and late fusion (vision) to enable robust multimodal reasoning and generation.
  • We establish an evaluation framework that systematically compares Cross-modal and Chain-of-Modality paradigms across consistent input-output modality configurations, while also assessing robustness to variations in accent and speaking speed.
@@ 2 Related Works
Speech-Enabled Vision-Language Models: LVLMs [27, 2, 3, 28] were initially designed for image-text reasoning, lacking native speech support. Early extensions added speech input and TTS output but remained limited in conversational expressiveness [29, 30], with higher latency and limited contextual adaptability due to multi-stage pipelines. GPT-4o [31]addresses these issues with native speech generation for real-time, expressive interaction, though its closed-source nature has driven open efforts to replicate its capabilities. Models like NExT-GPT [9], CoDi-1/2 [32, 33], Unified-IO [34], and AnyGPT [8] extend multimodal support via modality-specific encoders or projections, targeting perceptual tasks (e.g., image/audio/video generation, ASR, TTS). However, cross-modal reasoning (e.g., VQA, image captioning across text and speech modalities) remains underexplored. TMT [35] offers limited progress, enabling speech output from image captions but lacking joint multimodal reasoning.
Speech-Text-Vision Datasets: Most speech datasets used in multimodal learning—such as LibriSpeech [19], CommonVoice [20], and GigaSpeech [36]—are designed for ASR or TTS and do not include vision. While datasets like SpeechCOCO [37] and SpokenCOCO  [38] add speech to image captioning, they remain limited in scale, task diversity, and interactivity. Others, like How2 [39], offer aligned speech and video, but focus on narrow instructional domains and do not support flexible input-output modality switching. Some recent works synthesize multimodal dialogues by using large language models to generate text-based prompts, which are paired with audio and visual content using tools like TTS and image generation models. While useful for data augmentation, these datasets often focus on perception tasks and lack cognitively demanding reasoning challenges such as VQA or multimodal instruction following.
Direct vs. Instruction-Based Speech Generation: Speech generation in multi-modal systems generally follows one of two paradigms: direct generation, where speech is produced end-to-end without textual intermediates, and instruction-based generation, where models first generate text and then synthesize speech from it. Instruction-based approaches—used in models like GPT-4o [31], AnyGPT [8], NExT-GPT [9], and SpeechGPT [7]—offer modality consistency but often produce redundant or unnatural phrasing. In contrast, direct approaches aim to generate speech directly from inputs (e.g., Spectron [14]), allowing for more fluid prosody and conversational tone.
@@ 3 Data Generation
To construct a large-scale tri-modal dataset, we extend existing image-text corpora [24, 16] with corresponding spoken utterances. These datasets are selected for their abundance, diversity, and broad coverage of real-world visual and linguistic contexts, providing a strong foundation for scalable multimodal learning. We synthesize speech from text using the Melon-TTS model [40], chosen for its controllability over prosody and speaker characteristics. To enhance diversity, we vary speaking rates from 0.7× to 1.3× and include multiple English accents, such as American, British, Indian, and Australian. Additionally, we introduce environmental background noise—e.g., rain, footsteps, and ambient sounds—from the MUSAN corpus [25] to simulate realistic acoustic conditions. All audio is standardized to a 16 kHz sampling rate to ensure clarity and computational efficiency. This pipeline yields a rich and varied speech-text-image dataset suitable for both pre-training and fine-tuning multimodal models.
@@ 3.1 Pre-train dataset
Our pre-training strategy targets a unified speech-vision-language model capable of TTS, ASR, image captioning, and VQA. The training data includes: (1) 8M text-speech pairs from the publicly available LibriHeavy corpus [18] for speech generation and recognition, and (2) 6M image-text-speech triples, where speech is synthetically generated from image-caption pairs in LAION-COCO [16] and question-answer pairs from Visual Genome [17] to support vision-language reasoning. To enable modality switching, we use instruction-style prompts (e.g., “Answer this question in speech” or “Describe the image in text”), guiding the model to produce either spoken or written outputs. Speech outputs are capped at 5 seconds for training efficiency. Figure 1 shows the modality token distribution [^1] Token counts: text via Qwen-2.5-1.5B [41], speech via SpeechTokenizer [26], and images estimated at 256 tokens each.. Additional details are provided in Appendix A.
3.2 Visual Instruction dataset
For supervised fine-tuning (SFT), we adopt a similar structure to pre-training but introduce multi-turn conversations to improve coherence and long-range context retention. The text-speech subset includes LibriSpeech [19] and CommonVoice [20], covering diverse linguistic and command-oriented expressions. The vision-text-speech subset incorporates VQAv2, A-OKVQA [21], GQA [23], VizWiz [22], and COCO-Captions-2014, ensuring broad coverage of open-ended and visually grounded tasks. To enable consistent comparison across modalities, we design all tasks to support four input-output configurations: text-to-text, text-to-speech, speech-to-text, and speech-to-speech. Dialogue examples are constructed to fit within a 10-second speech limit and are paired with instruction prompts that guide the model to transition between text, speech, and vision as required. Additional details on the SFT dataset construction are provided in Appendix A.
4 Model Architecture
Our architecture is a hybrid of LLaVA [3] and AnyGPT [8], combining visual grounding with speech processing via discrete tokens. It integrates speech and vision inputs into a unified language model, denoted as $f_{\theta}(x)$, where $\theta$ represents model parameters. The overall architecture is shown in Figure 2.
@@ 4.3 Multimodal Fusion
Given text tokens $Z_{t}$, speech tokens $Z_{s}$, and visual embeddings $H_{v}$, we construct the multimodal input sequence:
$X=[\texttt{bos},\,\texttt{boi},\,\texttt{img},\,\texttt{eoi},\,Z_{t},\,\texttt{boa},\,Z_{s},\,\texttt{eoa},\,\texttt{eos}],$
where $n=|X|$ is the sequence length. Special tokens mark modality boundaries and control flow.
After embedding, we replace the placeholder token img with $H_{v}$, yielding:
$E^{\prime}=[$ $E_{\texttt{bos}},\,E_{\texttt{boi}},\,H_{v},\,E_{\texttt{eoi}},\,E_{Z_{t}},$
$E_{\texttt{boa}},\,E_{Z_{s}},\,E_{\texttt{eoa}},\,E_{\texttt{eos}}],$
with $E^{\prime}\in\mathbb{R}^{n^{\prime}\times d_{h}}$, where $n^{\prime}=n_{v}+n-1$. The LLM then processes $E^{\prime}$ for multimodal understanding.
@@ 4.4 Speech Decoding
4.5 Instruction-Based Speech Generation
During pre-training (see Appendix A.2), the model learns to generate both text and speech tokens directly. In SFT, text responses are always generated directly, while speech outputs follow two alternative strategies: direct generation or instruction-based generation, which we compare experimentally. For simpler tasks like ASR and TTS, we apply lightweight prompts (e.g., “The transcript of the given speech is:” for ASR, and “This is how your text sounds in speech:” for TTS). For more complex vision-language tasks—such as image captioning and VQA—we adopt an instruction-based approach: the model first generates a textual response, which is then converted into speech using structured prompts like “The textual caption is ‘caption’. Therefore, the audio caption is:”.
5 Experiments
5.1 Implementation Details
We use Qwen2.5-1.5B as our backbone LLM and train it with PyTorch. The vision encoder is CLIP-Large-Patch14-336 [42], which produces 256 tokens per image. Speech data is handled by the SpeechTokenizer [26], which encodes each 1-second segment of audio into 50 discrete tokens. When higher speech fidelity is required, SoundStorm-SpeechTokenizer [^2] https://github.com/ZhangXInFD/soundstorm-speechtokenizer extends this quantization approach to more nuanced tasks. For speech-input settings in image captioning and VQA, we use Melon-TTS to generate spoken questions or prompts, using the same configuration as in the training set, including a mix of speaking speeds and accents. More complementary details are provided in Appendix B.
5.2 Metrics
Text-Output tasks:
Speech-output tasks:
We evaluate speech generation using two methods: We transcribe model-generated speech with Whisper Medium [44] and compare the resulting text to human references (as in text-output tasks). However, ASR models can exhibit biases that introduce transcription errors, thereby distorting the perceived quality of the generated speech. Therefore, we use WavLM-TDNN [^3] https://github.com/yangdongchao/UniAudio/blob/main/UniAudio/tools/evaluation/compute_similarity_vc.pyto extract speech embeddings from both generated and reference speech and measure their similarity (via cosine similarity). This directly compares acoustic properties without relying on ASR.
5.3 Results
Model Backbone ASR TTS
LibriSpeech VCTK
WER WER Similarity
Human-level - 5.8 1.9 0.93
Wav2vec 2.0 [45] - 2.7 - -
Whisper Large V3 [44] - 1.8 - -
VALL-E [46] - - 7.9 0.75
VILA [47] Mixtral 8x7B [48] 8.1 - -
AnyGPT-7B LLaMA-2-7B 8.5 8.5 0.77
MIO-Ins [49] Yi-6B-Base [50] 10.3 4.2 -
Qwen2.5-Omni-7B [10] Qwen2.5-7B [50] 1.8 - -
SVLA-2B Qwen-1.5B 10.2 21.7 0.65
SVLA-2B-Text-Ins Qwen-1.5B 8.9 11.2 0.72
Text-Speech Performance: As shown in Table 1, SVLA-2B-Text-Ins outperforms SVLA-2B on both ASR and TTS, reducing ASR WER from 10.2 to 8.9 and TTS WER from 21.7 to 12 .2, while improving similarity from 0.65 to 0.72. Prompting with structured text (e.g., “This is the transcript:”) boosts performance by providing clearer context. While both models lag behind specialized systems like Whisper and Wav2vec 2.0 (WER 2.7), SVLA-2B-Text-Ins is competitive with AnyGPT-7B. The results highlight a trade-off: direct speech generation is more natural, but instruction-based prompts yield greater clarity and accuracy.